40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15272     Entry time: Thu Mar 12 16:13:16 2020
Author: gautam 
Type: Update 
Category: PSL 
Subject: Temperature sensors connected to Acromag 

[jordan, gautam]

the long DB25 cable to connect the Acromag chassis to the temperature sensor interface box arrived. We laid it out today. This cable does the following:

  1. Supply the box with +/- 24 V DC from the chassis. The pin arrangement here is rather unfortunate (the +/-24 V DC and GND pins are in close proximity), so if you're not careful, you'll create a short as you plug this cable in (as we found out today). So the best practise is to power down the crate before connecting/disconnecting this cable. Jordan will label this accordingly tomorrow.
  2. Pipe the "TAMB" and "TCAV" signals, corresponding to temperature sensors mounted to the PSL table-top and reference cavity exterior respectively, to the Acromags. We found during some initial testing that the "TAMB" signal was reaching the DB25 connector on the Acromag chassis, but wasn't going to any ADC channel - this was rectified.

Both signals now show up in the EPICS channels, but are noisy - I suspect this is because the return pin of the Acromag is not shorted to ground (this is a problem I've seen on the bench before). We will rectify this tomorrow as well.

We took this opportunity to remvoe the bench supply and temporary Acromag crate (formerly known as c1psl2) from under the PSL table. While trying to find some space to store the bench supply, we came across a damaged Oscilloscope in the second "Electronics" cabinet along the Y-arm, see Attachment #1. 

After this work, I found that the IMC autolocker was reliably failing to run the mcup script at the stage where the FSS gains are ramped up to their final values. I was, however, able to smoothly transition to the low-noise locked state if I was manipulating the EPICS sliders by hand. So I added an extra 2 seconds of sleep time between the increasing of the VCO gain to the final value and the ramping of the FSS gains in the mcup script (where previously there was none). Now the autolocker performs reliably.

Attachment 1: P_20200317_153736_vHDR_On_2.jpg  1.785 MB  Uploaded Tue Mar 17 22:01:22 2020  | Hide | Hide all
P_20200317_153736_vHDR_On_2.jpg
ELOG V3.1.3-